OpenStudy (anonymous):

I can't solve this problem of rotational dynamics. A solid disc is driven by an electrical motor through a pulley. The mass of the solid disc is 80 kg and its radious 0.625 meters. The pulley is welded to the disc, having the same rotation axis. The rotational movement is transmitted to the disc-pulley system by a belt. The tension of the upper segment of the belt is 135 N, and the tension on the lower segment is unknown. How can I find this unknown tension? The angular acceleration of the disc-pulley is 1.67 radians per second squared.

OpenStudy (anonymous):

Here is a diagram of the system. I thought it could be solved using the equation \[\tau = I * \alpha \] (moment is equal to the moment of inertia times the angular acceleration). But the mass of the pulley is not given, then I cannot calculate total moment of inertia.

OpenStudy (anonymous):

Both the pulley and disc are "welded" =>Alpha of disc =Alpha of Pulley disc system. Its not given which direction the pulley is accelerating . So i am assuming T=135 is the greater tensile force =>135*R-T*R=I(disc)*alpha From this we can calculate T(Unknown Tensile force). If we assume T>135 then T*R-135*R=I(disc)*Alpha